A man with a height 1.8 meters walking towards a 46 meters height building with a 1.2 m/s velocity. You are standing on top of the building to throw an egg at the man's head. At what distance from the building should the man be, so if you drop an egg, you can hit his head?
1
Expert's answer
2017-09-14T07:05:06-0400
The time of the egg free falling from the top of the building to the man head is
t=√((2(H-h))/g).
So, distance from the building
l=vt=v√((2(H-h))/g)=1.2×√((2(46-1.8))/9.8)=3.6 m.
Answer: 3.6 m